# Triangular Automata The 256 Elementary Cellular Automata of the 2D Plane

### v5

@@ Abstract
Triangular Automata (TA) stands for cellular automata in the triangular grid. This work focuses on the simplest type of TA called Elementary Triangular Automata (ETA). They are argued to be the two-dimensional counterpart of Wolfram’s Elementary Cellular Automata. Conceptual and computational tools for their study are presented along with an initial analysis. The paper is accompanied by a website where the results can be explored interactively [^1] Website: paulcousin.github.io/triangular-automata. The source code is available in the form of a Mathematica package [^2] Package demonstration: paulcousin.github.io/triangular-automata-mathematica.
@@ 1 Introduction
Cellular automata in the triangular grid, or Triangular Automata (TA) for short, have already been studied in a few papers [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. This work will focus on a natural subset of TA called Elementary Triangular Automata (ETA).
@@
ETA cells hold only binary states, each cell will thus either be:
@@
  2. “dead” and colored white [Uncaptioned image] with a state $s=0$
ETA rules determine the future state of a cell based on its current state and the states of its neighbors, regardless of their orientation. This results in only 8 possible local configurations.
@@
This paper uses a graph-theoretical framework developed in a previous work on Graph-Rewriting Automata[18]. The triangular grid will here be considered as a graph (Figure 3). This graph must be expanded at each time step to simulate an infinite grid. The region of influence of a single cell grows in hexagonal layers (Figure 3(a)). This is therefore the most efficient way to expand the graph as well. How to do this in practice will be detailed in Section 3.2.
@@
It is useful to see the triangular grid as a graph because computing the evolution of ETA is made quite easy by properties of its adjacency matrix $\mathcal{A}$ and statevector $\mathcal{S}$. Every vertex $v$ of this graph will hold a state $s(v)$. The neighborhood $N(v)$ of a vertex is defined as the set of its adjacent vertices.
@@
The configuration $c(v)$ of a vertex is a number which, when they are indexed as in Figure 2, can be expressed as follows:
@@
The space of possible ETA rules is finite. For each of the 8 configurations, a rule must specify whether the vertex will be dead or alive at $t+1$. Consequently, there are only $2562^{8}=$ 256 possible rules. For this reason, ETA can be seen as the two-dimensional counterpart of Wolfram’s 256 Elementary Cellular Automata[19, 20]. In fact, there is a direct equivalence between 64 of each (see Section 2.6). Furthermore, the triangle is the regular polygon tiling 2D space with the smallest number of neighbors per cell. ETA are thus the most basic 2D cellular automata and have a fundamental aspect in this regard.
Each rule $R$ is a map from configuration space to state space.
@@
Each rule can be labeled by a unique rule number $n$ (see Equation 4). We will use the labeling system which was independently proposed in [8] and [18], since it must be somewhat natural and because it has useful properties. This system, inspired by the Wolfram code [19], is such that a rule number in its binary form displays the behavior of the rule. Starting from the right, its digits indicate the future state for each configuration as they have been ordered previously. Figure 4 shows the example of rule 210.
@@
Figure 4: Rule $210=1101,0010_{2}$
@@ 2 Behavior
Before going into the details of how to compute these automata, we can take a look at how they behave. In this section, a preliminary study of ETA will be presented to motivate a future, more in-depth analysis. It is of particular interest to see what happens to a single living cell under different ETA rules so, unless otherwise noted, the following figures come from this starting point.
@@ 2.3 Chaos
Given the existing literature on cellular automata, it is quite expected to see some of these rules behave chaotically. The example of rule 53 confirms it. Starting from two randomly generated 64 layers wide grids that are completely similar except for the central cell which is alive in (1) and dead in (2), the trajectories strongly diverge.
@@
Figure 8: Rule 53
@@ 2.4 Self-Reproduction
As mentioned in [17], one of the original motivations for the development of cellular automata was to create a mathematical model of self-reproduction. Interestingly, 4 of the 256 ETA rules naturally reproduce any finite pattern given as initial conditions: rules 85, 90, 165 and 170. A proof of self-reproduction based on path counting already exists for rule 170 [17]. Similarly spirited proofs could probably be proposed for the others.
{tblr}
—Q[c,m]—Q[c,h] Q[c,h] Q[c,h] Q[c,h]— t rule 85 rule 90 rule 165 rule 170
@@
Table 1: Pattern self-reproduction from a recognizable starting point
@@ 2.5 Space-Time
Similar to the way Elementary Cellular Automata [19, 20] are most often represented, the evolution of an ETA can be displayed in one single plot. Here, an instant is two-dimensional, so adding the dimension of time creates a 3D structure. In these space-time plots, time flows downward. The successive grids are stacked beneath each other, starting from the initial conditions at the top. To avoid the infinite planes created by an alive environment, we can display only the cells that have the opposite state to it at each time step. A lot of information is therefore lost. We do not see most of the internal structure and we cannot know the state of the environment. Nevertheless, this representation helps visualize some properties of ETA that are difficult to notice otherwise. For instance, certain rules create 3D space-time fractals.
@@
Another way to see the time evolution of an ETA in a single plot is to look at just one line of the grid, i.e. to look at a sliced space-time plot from the side. There is a most natural way to slice: along the lines that preserve the neighborhood relations.
@@
The states along these lines can be plotted beneath each other to create Wolfram-style plots, here called slice plots.
@@ 2.6 Elementary Cellular Automata
Under a certain condition, ETA rules maintain the uniformity of layers. This is the case when the $2^{\text{nd}}$ and $3^{\text{rd}}$ digits of the binary rule number are the same, as well as the $6^{\text{th}}$ and $7^{\text{th}}$.
@@
64 ETA rules satisfy this condition and they are all striclty equivalent to a unique symmetric Elementary Cellular Automaton (ECA) [19].
@@ 2.7 Center Column
Figure 15: Center column $1\leq t\leq 4096$ (read left-to-right, top-to-bottom)
The symmetries of the triangle impose that, starting from a single living cell, layers 1 and 2 remain uniform. We therefore have a three layers wide pillar, which is fully described by a sequence of bit triplets. These can be mapped to RGB colors to extend the previous plots.
@@
Figure 18: Three center columns $1\leq t\leq 4096$
@@ 2.8 Rule 210
Rule 210 is easy to describe: “change the state of cells that have one neighbor alive”. It is special for several reasons: it creates larger structures as time goes on and a new type of structure appears from around $t=1024$.
{tblr}
c c c c c Refer to caption & Refer to caption Refer to caption Refer to caption Refer to caption
@@
Rule 210 initially alternates between what seems like 2 population densities (i.e. number of living cells / size of the first cell’s region of influence). When the new structure appears, around $t=1024$, the population stabilizes in a density cycle of period 4 between slowly evolving values.
@@ 2.10 Textures
(a) Rule 100 at at $t=64$
(b) Rule 108 at at $t=512$
(c) Rule 204 at at $t=32$
@@ 2.11 Boring Rules
There is an identity rule which leaves any grid unchanged: rule 240.
@@
And a negative rule that swaps alive and dead states: rule 15.
@@ 2.12 Twins
A simple procedure can be followed to find the evil twin of a rule that has the same effect but in the negative world. To find it, take the number in its binary form (with the leading zeros needed for the number to be 8 digits long), swap ones and zeros and read it backwards. Let us take rule 214 as an example.
@@ 3.1 Tools
    $\diamond$ here represents an operator that joins matrices corner to corner.
@@
    $\begin{pmatrix}a&b&c&{\color[rgb]{1,.5,0}d}\\ e&f&g&{\color[rgb]{1,.5,0}h}\\ {\color[rgb]{1,0,0}i}&{\color[rgb]{1,0,0}j}&{\color[rgb]{1,0,0}k}&{\color[rgb]{.75,0,.25}l}\end{pmatrix}^{\searrow}=\begin{pmatrix}{\color[rgb]{.75,0,.25}l}&{\color[rgb]{1,0,0}i}&{\color[rgb]{1,0,0}j}&{\color[rgb]{1,0,0}k}\\ {\color[rgb]{1,.5,0}d}&a&b&c\\ {\color[rgb]{1,.5,0}h}&e&f&g\\ \end{pmatrix}$ (7)
@@
    $\mathbb{I}_{i}$ is the $i\times i$ identity matrix. $\mathbb{S}_{i}$ will be a $i\times(i+1)$ “stairs” matrix.
@@ 3.2 Growing the triangular grid
As mentioned in Section 1, the grid will be grown from a single cell by adding layers. The cells will be ordered counterclockwise, with the first vertex of each new layer placed on the south-east diagonal (see Figure 29).
Figure 29: Cell ordering
This is done with a precursor of the adjacency matrix called the grid matrix $\mathcal{G}$. Up to the third layer, the grid matrix is better hand-coded. The third matrix $\mathcal{G}_{3}$ can be seen in Figure 30. The subsequent layers will follow a repeating pattern.
@@
Each layer consists in a sub-matrix and the grid matrix will be :
@@
From $m_{4}$, these sub-matrices become:
$m_{i}=\begin{cases}\mathbb{S}_{\frac{i}{2}}\diamond\mathbb{I}_{(\frac{i}{2}-1)}\diamond\mathbb{S}_{\frac{i}{2}}\diamond\mathbb{I}_{(\frac{i}{2}-1)}\diamond\mathbb{S}_{\frac{i}{2}}\diamond\mathbb{I}_{(\frac{i}{2}-1)}&\text{ if $i$ is even}\\ \vspace{-2mm}\\ \left(\mathbb{I}_{\lceil\frac{i}{2}-2\rceil}\diamond\mathbb{S}_{\lceil\frac{i}{2}\rceil}\diamond\mathbb{I}_{\lceil\frac{i}{2}-2\rceil}\diamond\mathbb{S}_{\lceil\frac{i}{2}\rceil}\diamond\mathbb{I}_{\lceil\frac{i}{2}-2\rceil}\diamond\mathbb{S}_{\lceil\frac{i}{2}\rceil}\right)^{\searrow}&\text{ if $i$ is odd}\end{cases}$ (12)
Assuming that $\mathbb{I}_{0}$ is a $0\times 0$ matrix, this pattern actually holds for $m_{2}$ and $m_{3}$. However, depending on how this is implemented, coding the first 3 layers by hand might be the best option.
@@ 3.3 Evolving the state
The environment will be simulated by using two layers around the region of the influence of our initial structure. If the initial structure is a single triangle, then the computed grid will contain $t+2$ layers.
@@
    First, a layer is added with the same state as the last vertex (both the grid/adjacency matrix and the state vector must be updated).
@@
    Second, a configuration vector $\mathcal{C}$ is computed ($o$ is the order of the graph).
@@
    Finally, the state of all vertices of the last layer (created in step 1) is set to the value of the last vertex of the now penultimate layer. This removes the artefacts coming from the edges of the computed grid.
@@
  1. Evolving the state of the grid is where this framework pays off the most. Steps 2 and 3 are mathematically sufficient if we consider working in an infinite graph. They would also be the only steps required in a closed grid, like a triangulated surface [8] for example.
@@
  3. Step 1 can be avoided in the case where the three outermost layers have a uniform state, which is easy to check.
@@ 3.4 Plotting the result
{tblr}
[…]
@@
6:$\mathcal{K}.append\Big{(}\mathcal{K}\big{[}-3(l-1)\big{]}+step\Big{)}$triangleright $\mathcal{K}[-n]\equiv n^{th}$ coords from the end
7:for $i\leftarrow 0,3l-2$ do
@@ 4 Conclusion
The triangular tessellation plays an important role in many disciplines, from computer graphics to architecture. TA are a way of populating it with aesthetic patterns. Beyond the possible applications, ETA are somewhat fundamental cellular automata, making them an elegant model of complexity. With the framework presented here, the 256 ETA rules can now be thoroughly explored, even with limited computational resources.
@@
    ETA rules could be classified according to some common criteria,
@@
    the approach taken in this paper could easily be applied to a wider class of TA,
@@
    a GPU-accelerated implementation could be used to explore longer timescales, as has already been done for Graph-Rewriting Automata[21],

### v7

@@ Abstract
Cellular automata on the triangular grid are here referred to asTriangular Automata (TA). This paper focuses on the simplest class of TA, called Elementary Triangular Automata (ETA). They are argued to be the two-dimensional counterpart of Wolfram’s Elementary Cellular Automata. Conceptual and computational tools for their study are presented, along with an initial analysis. This paper is accompanied by a website [^1] Website: https://triangular-automata.net where the results can be explored interactively. The source code is available in the form of a Mathematica package [^2] Package demonstration: https://triangular-automata.net/code.
@@ 1 Introduction
Cellular automata on the triangular grid (Figure 1), or Triangular Automata (TA) for short, have already been studied in a few papers [gerlingClassificationTriangularHoneycomb1990, baysCellularAutomataTriangular1994, imaiComputationuniversalTwodimensional8state2000, naumovGeneralizedCoordinatesCellular2003, linApplicationUnstructuredCellular2009, baysCellularAutomataTriangular2009, baysGameLifeNonsquare2010, zawidzkiApplicationSemitotalistic2D2011, brecklingCellularAutomataEcological2011, saadatCellularAutomataTriangular2016, ortigozaACFUEGOSUnstructuredTriangular2016, uguzStructureReversibility2D2017, saadatCellularAutomataApproach2018, wainerIntroductionCellularAutomata2019, pavlovaUsingCellularAutomata2020, saadatGeneratingPatternsTriangular2021, saadatCopyMachinesSelfreproduction2023]. This paper focuses on a natural subset of TA called Elementary TA (ETA).
@@
ETA cells hold only binary states; each cell can thus be either:
@@
  2. “dead” and colored white [Uncaptioned image], with a state $s=0$
ETA rules determine the future state of a cell based on its current state and the states of its neighbors, regardless of their orientation. This results in only 8 possible local configurations, as shown in Figure 2.
@@
This paper uses a graph-theoretical framework developed in a previous work on Graph-Rewriting Automata [cousin2022organic]. The triangular grid will here be considered as a graph (Figure 3). This graph must be expanded at each time step to simulate an infinite grid. The region of influence of a single cell grows in hexagonal layers (Figure 3(a)). This is therefore the most efficient way to expand the graph as well. How to do this in practice is detailed in Section 3.2.
@@
It is useful to see the triangular grid as a graph because computing the evolution of an ETA is made relatively straightforward by properties of its adjacency matrix $\mathcal{A}$and state vector $\mathcal{S}$. Every vertex $v$ of this graph will hold a state $s(v)$. The neighborhood $N(v)$ of a vertex is defined as the set of its adjacent vertices.
@@
The configuration $c(v)$ of a vertex is a number that, when indexed as in Figure 2,can be expressed as:
@@
The space of possible ETA rules is finite. For each of the 8 configurations, a rule must specify whether the vertex will be dead or alive at $t+1$. Consequently, there are only $2^{8}=256$ possible rules. For this reason, ETA can be seen as the two-dimensional counterpart of Wolfram’s 256 Elementary Cellular Automata [wolfram2002new, weisstein2002elementary]. In fact, there is a direct equivalence between 64 of each (see Section 2.7). Furthermore, the triangle is the regular polygon that tiles the 2D plane with the smallest number of neighbors per cell. ETA are thus the most basic two-dimensional cellular automata and are fundamental in this regard.
A rule $R$ is a map from configuration space to state space.
@@
Each rule can be labeled by a unique rule number $n$:
@@
This labeling system was proposed independently in [zawidzkiApplicationSemitotalistic2D2011] and [cousin2022organic] and will be used in this paper, since it is fairly natural and has useful properties. This system, inspired by the Wolfram code [wolfram2002new], is such that a rule number in its binary form displays the behavior of the rule. Starting from the right, its digits indicate the future state for each configuration as they have been ordered previously. Figure 4 shows the example of rule 210.
Figure 4: Rule $210=1101\,0010_{2}$
@@ 2 Behavior
Before going into the details of how to compute these automata, we can take a look at how they behave. In this section, a preliminary study of ETA is presented to motivate a future, more in-depth analysis. It is of particular interest to see what happens to a single living cell under different ETA rules, so, unless otherwise noted, the following figures come from this starting point.
@@ 2.3 Chaos
Given the existing literature on cellular automata, it is not surprising that some rules behave chaotically. The example of rule 53 confirms it. Starting from two randomly generated 64-layer-wide grids that are completely similar except for the central cell which is alive in (1) and dead in (2), the trajectories strongly diverge.
@@
2.4 Gliders
Figures 8 and 9 show two small gliders.
Figure 8: A glider of rule 21
Figure 9: A glider of rule 53
@@ 2.5 Self-Reproduction
As mentioned in [saadatCopyMachinesSelfreproduction2023], one of the original motivations for the development of cellular automata was to create a mathematical model of self-reproduction. Interestingly, 4 ofthe 256 ETA rules naturally replicate any finite pattern given as initial conditions: rules 85, 90, 165 and 170. A proof of self-reproduction based on path counting already exists for rule 170 [saadatCopyMachinesSelfreproduction2023]. Similar proofs could probably be proposed for the others.
t rule 85 rule 90 rule 165 rule 170
@@
Table 1: Pattern replication from a recognizable starting point
@@ 2.6 Space-Time
Similar to the way Elementary Cellular Automata [wolfram2002new, weisstein2002elementary] are most often represented, the evolution of an ETA can be displayed in a single plot. Here, an instant is two-dimensional, so adding the dimension of time creates a 3D structure. In these space-time plots, time flows downward. The successive grids are stacked beneath each other, starting from the initial conditions at the top. When the background state is 1,we can display the dead cells to avoid the infinite planes. A lot of information is therefore lost. We do not see most of the internal structure and we cannot know the state of the universe. Nevertheless, this representation helps visualize some properties of ETA that are difficult to notice otherwise. For instance, certain rules create 3D space-time fractals, as in Figure 10.
@@
Another way to see the time evolution of an ETA in a single plot is to look at just one line of the grid, that is, to look at a sliced space-time plot from the side. There is a most natural way to slice: the lines parallel to the sides of the triangles preserve the neighborhood relations (Figure 11).
@@
The states along these lines can be plotted beneath each other to create Wolfram-style plots, here called slice plots (Figure 12).
@@ 2.7 Elementary Cellular Automata
Under a certain condition, ETA rules maintain the uniformity of layers. This is the case when the second and third digits of the binary rule number are the same, as well as the sixth and seventh.
@@
64 ETA rules satisfy this criterion. When the initial conditions have uniform layers (e.g., a single living cell or infinite stripes), these rules are each strictly equivalent to a unique symmetric Elementary Cellular Automaton (ECA) [wolfram2002new].
@@ 2.8 Center Column
Figure 16: Center column for $1\leq t\leq 4096$ (read left-to-right, top-to-bottom)
The symmetries of the triangle impose that, starting from a single living cell, layers 1 and 2 remain uniform. We therefore have a three-layer-wide pillar, which is fully described by a sequence of bit triplets. These can be mapped to RGB colors (Figures 17and 18) to extend the previous plots.
@@
Figure 19: Three center columns, $1\leq t\leq 4096$
@@ 2.9 Rule 210
Rule 210 is easy to describe: “cells with one living neighbor change state”. It is special for several reasons: it creates larger structures as time goes on, and a new type of structure with very large features appears from around $t=1024$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Rule 210 initially alternates between what seems like 2 population densities (i.e., number of living cells / size of the first cell’s region of influence). When the new structure appears, around $t=1024$, the population stabilizes in a density cycle of period 4 between slowly evolving values.
@@ 2.11 Textures
(a) Rule 100 at $t=64$
(b) Rule 108 at $t=512$
(c) Rule 204 at $t=32$
@@ 2.12 Boring Rules
Rule 240 is the identity and leaves any grid unchanged. Rule 15 is the negative rule that swaps alive and dead states. Rule 0 kills everything, and rule 255 makes all cells live.
@@ 2.13 Twins
A simple procedure can be followed to find the evil twin of a rule that has the same effect but in the negative world. To find it, take the number in its binary form (with the leading zeros needed to make it 8 digits long), swap ones and zeros and read it backward. Let us take rule 214 as an example.
@@ 3.1 Tools
    $\diamond$ here represents an operator joining matrices corner to corner.
@@
    $\begin{pmatrix}a&b&c&{\color[rgb]{1,.5,0}d}\\ e&f&g&{\color[rgb]{1,.5,0}h}\\ {\color[rgb]{1,0,0}i}&{\color[rgb]{1,0,0}j}&{\color[rgb]{1,0,0}k}&{\color[rgb]{.75,0,.25}l}\end{pmatrix}^{\hskip-4.2679pt\searrow}=\begin{pmatrix}{\color[rgb]{.75,0,.25}l}&{\color[rgb]{1,0,0}i}&{\color[rgb]{1,0,0}j}&{\color[rgb]{1,0,0}k}\\ {\color[rgb]{1,.5,0}d}&a&b&c\\ {\color[rgb]{1,.5,0}h}&e&f&g\\ \end{pmatrix}$ (7)
@@
    $\mathbb{I}_{i}$ is the $i\times i$ identity matrix. $\mathbb{S}_{i}$ will be an $i\times(i+1)$ “stairs” matrix.
@@ 3.2 Growing the triangular grid
As mentioned in Section 1, the grid will be grown from a single cell by adding layers. The cells will be ordered counterclockwise, with the first vertex of each new layer placed on the southeast diagonal (see Figure 30).
Figure 30: Ordering of cells
This is done by using a precursor of the adjacency matrix called the grid matrix $\mathcal{G}$.For the first three layers, the grid matrix is best hand-coded. The third matrix $\mathcal{G}_{3}$ can be seen in Figure 31. The subsequent layers will follow a repeating pattern.
@@
Each layer consists of a submatrix and the grid matrix will be :
@@
From $m_{4}$, these submatrices become:
[…]
Assuming that $\mathbb{I}_{0}$ is a $0\times 0$ matrix, this pattern actually holds for $m_{2}$ and $m_{3}$. However, depending on how this is implemented, inputting the first 3 layers by hand might be the best option.
@@ 3.3 Evolving the state
The universe will be simulated using two layers around the region of influence of our initial structure. If the initial structure is a single triangle, then the computed grid will contain $t+2$ layers.
@@
    A layer is added with the same state as the last vertex (both the grid/adjacency matrix and the state vector must be updated).
@@
    A configuration vector $\mathcal{C}$ is computed ($o$ is the order of the graph).
@@
    The state of all vertices of the last layer (created in step 1) is set to the value of the last vertex of the now penultimate layer. This removes the artifacts coming from the edges of the computed grid.
@@
  1. Evolving the state of the grid is where this framework pays off the most. Steps 2 and 3 are mathematically sufficient if we consider working in an infinite graph. They would also be the only steps required in a closed grid, like a triangulated surface [zawidzkiApplicationSemitotalistic2D2011], for example.
@@
  3. Step 1 can be avoided if the three outermost layers have a uniform state, which is easy to check.
@@ 3.4 Plotting the result
Layer Coordinates of the base triangle Illustration
@@
[…]
7:for i=0,3l2i=0,3l-2 do
@@ 4 Conclusion
The triangular tessellation plays an important role in many disciplines, from computer graphics to architecture. Triangular Automata (TA) are a way of populating the triangular grid with aesthetic patterns. Beyond the possible applications, Elementary Triangular Automata (ETA) are somewhat fundamental cellular automata, making them an elegant model of complexity. With the framework presented here, the 256 ETA rules can now be thoroughly explored, even with limited computational resources.
@@
    ETA rules could be classified according to some common criteria.
@@
    The approach taken in this paper could easily be applied to a wider class of TA.
@@
    A GPU-accelerated implementation could be used to explore longer timescales, as has already been done for Graph-Rewriting Automata [GRAweb].